OpenStudy (anonymous):

A cone-shaped hole is drilled into a solid cube of metal. If the cube has side length 9 cm, what is the volume of the metal after the hole is drilled? Use 3.14 and round your answer to the nearest tenth.

OpenStudy (anonymous):

assuming the radius of the cone = 4.5 and the height = 9

OpenStudy (anonymous):

that is you are trying to get the largest cone shape

OpenStudy (anonymous):

Volume of cone = (1/3) pi r^2 h Volume after drilling = h^3 - (1/3) pi r^2 h

OpenStudy (anonymous):

\[V = h ( h^2 - \frac{1}{3} \pi r^2 ) \]

OpenStudy (anonymous):

sub in values r=4.5 , h=9